### v1

@@ 4.3 Comparison with State-of-the-Arts
Table 3: Cross-manipulation detection AUC on the three unseen manipulation methods (C23).
@@
Figure 4: Robustness to various common perturbations at five severity levels.
Robustness to Real-World Perturbations. Images and videos transmitted online always undergo various perturbations that erase forgery cues within the image/video contents. We introduce common perturbations to measure the model’s robustness: Gaussian blurring, pink noise, white noise, and blockwise. Each perturbation type includes five severity levels to simulate diverse real-world conditions. As shown in Fig. 4, the detectors’ performance consistently deteriorates with increasing severity levels. However, our proposed MoE-FFD exhibits significantly greater resilience to most perturbations compared to previous methods. Notably, MoE-FFD achieves substantial improvements in robustness compared to the ViT-B baseline.
Discussion. MoE-FFD achieves superior generalizability and robustness compared to previous methods, which can be attributed to the following designs: (1). MoE-FFD only updates external modules while preserving the abundant ImageNet knowledge, enabling the model to adaptively learn forgery-specific features; (2). MoE-FFD integrates the LoRA and Convpass Adapter with the ViT backbone, effectively leveraging the expressivity of transformers and the local forgery priors of CNNs; (3). The incorporation of MoE modules facilitates optimal selection of LoRA and Adapter experts for forgery feature mining. Furthermore, MoE-FFD presents a parameter-efficient approach to face forgery detection due to the utilized PEFT strategy. To validate the effectiveness of the designed components, we perform ablation experiments next.
Table 4: Ablation study on the designed PEFT modules.
ViT-B MoE MoE CDF DFDC WDF
@@
Table 5: Cross-dataset detection performance on different backbones.
Backbone MoE-FFD #Params CDF DFDC WDF
@@
Table 6: Impacts of the matrix rank of LoRA layers and the effectiveness of MoE.
Setting CDF DFDC WDF
@@
Table 7: Impacts of different designed Adapters and the effectiveness of MoE.
Setting CDF DFDC WDF
@@
4.4 Ablation Experiments
4.4.1 Impacts of the PEFT Layers.
MoE-FFD integrates one LoRA layer and one Adapter layer with each ViT block. The LoRA layer is designed to learn forgery-specific parameters for subsequent attention mechanisms, capturing long-range interactions within the input faces. Meanwhile, the Convpass Adapter introduces forgery local priors into the plain ViT model. To study the effectiveness of the designed PEFT layers, we report the cross-dataset detection results in Table 4.3. Utilizing MoE LoRA and MoE Adapter significantly improves detection performance across all datasets compared to the ViT-B baseline. This enhancement is attributed to the LoRA and Adapter layers effectively retaining the ImageNet knowledge while adaptively learning forgery features. Furthermore, the combination of LoRA and Adapter layers allows the model to capture both long-range interactions and local forgery cues, further enhancing its performance.
4.4.2 Effectiveness on Other Backbones.
Our proposed method can be flexibly adapted to other vision transformer backbones in a plug-and-play manner. To validate this adaptability, we incorporate MoE-FFD into ViT-Tiny, ViT-Small, and ViT-Large models. The cross-dataset results are presented in Table 4.3. By using MoE-FFD, significant performance improvements are observed across the ViT backbones. This, in turn, demonstrates the flexibility and effectiveness of our method.
Figure 5: LoRA expert selection frequency on (a). CDF and (b). WDF datasets; Adapter expert selection frequency on (c). CDF and (d). WDF datasets.
Table 8: MoE v.s. Multi-Experts.
Method Training Inference CDF DFDC WDF
@@ 4.4 Ablation Experiments
Table 9: Effectiveness of loss components.
$L_{ce}$ $L_{moe}$ CDF DFDC WDF
AUC EER AUC EER AUC EER
- 82.65 25.57 78.14 29.06 76.29 30.43
86.69 22.06 80.83 26.67 80.64 27.11
4.4.3 Impacts of the MoE Learning Scheme.
The proposed MoE is designed to dynamically select the optimal experts for face forgery detection. In Table 4.3, we compare the performance of the proposed MoE with individual LoRA experts of varying ranks. Results show that MoE consistently outperforms individual LoRA experts on all three datasets. Notably, different datasets exhibit varying optimal LoRA ranks. Similarly, we investigate the impact of different Adapters in Table 4.3. Adapters tend to capture different local features of input faces. MoE again achieves the superior results compared to using a single Adapter. It should be noted that the MoE approach only introduces negligible additional activated parameters, demonstrating that the performance improvements stem from the proposed MoE learning scheme.
In Fig. 5, we further present the expert selection distributions on two datasets. Fig. 5 (a) and (b) show the LoRA expert selection frequency on CDF and WDF datasets, while Fig. 5 (c) and (d) illustrate the Adapter expert selection frequency. Different Deepfake datasets generally exhibit dramatic domain gaps. From Fig. 5, we observe distinct expert selection distributions between the CDF and WDF datasets. This finding demonstrates that MoE-FFD effectively selects the optimal experts for different data from another perspective.
@@
To further demonstrate the dynamic selection of optimal LoRA and Adapter experts by MoE, we compare MoE with Multi-Experts (Multi-E). Multi-E aggregates the features of all designed experts. The efficiency and the face forgery detection performance of MoE and Multi-E are reported in Table 4.4.2. Results indicate that MoE achieves a 1.28$\times$ speedup in training and 1.23$\times$ speedup in inference. Despite these efficiency gains, MoE consistently outperforms Multi-E in terms of detection performance. This observation further underscores the effectiveness of the proposed method in selecting the optimal expert for face forgery detection.
Figure 6: T-SNE feature distributions of the ViT baseline and MoE-FFD. Visualization results of (a). ViT baseline on CDF dataset, (b). MoE-FFD on CDF dataset, (c). ViT baseline on WDF dataset, and (d). MoE-FFD on WDF dataset.
Figure 7: Grad-CAM maps of the baseline model (ViT) and our proposed method MoE-FFD on four Deepfake datasets: FF++/DF, CDF, DFDC, and WDF.
@@
The gating network often exhibits a tendency to consistently assign large weights to only few experts [56]. To address this issue, we introduce an MoE loss component aimed at encouraging equal importance among all experts. The results in Table 4.4.2 demonstrate that $L_{moe}$ effectively mitigates the gate overfitting problem and boosts the detection performance.
@@ 4.5 Visualization Results
We further provide the Grad-CAM maps of the baseline model and MoE-FFD in Fig. 7. Both models are trained on FF++ dataset and tested on four Deepfake datasets: (a). FF++/DF, (b). CDF, (c). DFDC, and (d). WDF. While the baseline model often neglects informative fake facial regions or attends to peripheral irrelevant areas, MoE-FFD consistently directs attention to manipulated regions within input face queries. This observation highlights the superiority of our method.
5 Conclusions
In this paper, we introduced MoE-FFD, a generalized yet parameter-efficient method for detecting face forgeries. By incorporating external lightweight LoRA and Adapter layers with the frozen ViT backbone, our framework adaptly acquired forgery-specific knowledge with minimal activated parameters. This approach not only harnesses the expressiveness of transformers but also capitalizes on the local forgery priors in CNNs, contributing to enhanced detection performance. Through dynamic expert selection within both LoRA and Adapter layers, our MoE design further enhances the model’s generalizability and robustness. Extensive experiments consistently demonstrated MoE-FFD’s superiority in face forgery detection across diverse datasets, manipulation types, and perturbation scenarios. Moreover, MoE-FFD serves as a parameter-efficient detector and can seamlessly adapt to various ViT backbones, facilitating its deployment and fine-tuning in real-world applications. Last but not least, comprehensive ablation studies demonstrated the effectiveness of our designed LoRA layers and Convpass Adapter layers, and the MoE modules indeed helped the model to search the optimal forgery experts. We anticipate that MoE-FFD will inspire future advancements in face forgery detection, particularly in bolstering generalization and efficiency.

### v2

@@ 4.1 Implementation Details
4.2 Datasets and Evaluation Metrics
The experiments encompass the following six datasets for training and testing: FaceForensics++ (FF++) [72], CelebDF-v2 (CDF) [27], WildDeepfake (WDF) [73], DeepFake Detection Challenge Preview (DFDC-P) [74], DeepFakeDetection (DFD) [75], and DeepForensics-1.0 (DFR) [76]. FF++ is a widely used dataset in face forgery detection, which includes four face manipulation types: Deepfakes (DF) [77], Face2Face (FF) [78], FaceSwap (FS) [79], and NeuralTextures (NT) [80]. The remaining datasets represent recent five Deepfake datasets with diverse environment variables and various video qualities. We perform cross-dataset and cross-manipulation evaluations to examine the models’ generalization capability. Additionally, we measure the model’s robustness by applying it to perturbed face images with various perturbation types and severity levels. Consistent with prior arts, we adopt Area under the ROC Curve (AUC) and Equal Error Rate (EER) as evaluation metrics. AUC measures the Area under the Receiver Operating Characteristic (ROC) curve, while EER denotes the False Positive Rate (FPR) that equals to the True Positive Rate (TPR).
TABLE III: Cross-manipulation detection AUC on the remaining three unseen manipulation techniques (FF++ C23 dataset).
@@ 4.2 Datasets and Evaluation Metrics
4.3 Comparison with State-of-the-Arts
4.3.1 Cross-Dataset Evaluations.
In this subsection, we perform cross-dataset evaluations to evaluate the models’ generalization capability. In practical scenarios, the trained models are vulnerable to unseen domains, leading to significant performance degradations. We train all face forgery detectors on the FF++ (C23) dataset and directly evaluate them on five unseen Deepfake datasets: CDF, WDF, DFDC-P, DFD, and DFR. Table I presents the frame-level and video-level detection results, divided into top and bottom sections. Video-level detection results highlighted in blue color are obtained by averaging all frame scores within each video. We bold the best results and underline the second-best results. Notably, MoE-FFD achieves the highest detection performance on four datasets in both frame-level and video-level detection. Additionally, MoE-FFD’s performance is on par with CFM on the DFD dataset. Moreover, MoE-FFD outperforms CFM by a clear margin in average performance. The average frame-level AUC improves by 2.86%, rising from 83.94% to 86.78%. Similarly, MoE-FFD also outperforms others in video-level detection on average. Our MoE-FFD’s generalizability improvements over existing methods offers valuable insights into the effectiveness of the proposed approach.
Figure 4: Robustness to various common perturbations at five severity levels: Gaussian blur, pink noise, white noise, and block wise.
4.3.2 Cross-Manipulation Evaluations.
Existing face forgery detectors often struggle to handle emerging manipulation techniques. With the rapid development of AIGC, more sophisticated manipulation techniques continue to emerge, posing challenges for existing detection models. Ensuring model generalizability to unseen forgeries is crucial for real-world applications. In this study, we conduct cross-manipulation experiments involving four forgery techniques: Deepfakes (DF), Face2Face (FF), FaceSwap (FS), and NeuralTextures (NT). Table II presents the results, where models trained on three manipulation types are tested on the remaining one. MoE-FFD exhibits the best average detection results. Compared to the ViT-B baseline, the proposed method achieves an average AUC enhancement of 18.5%, going from 62.3% to 80.8%. In Table III, we further examine models trained on one manipulation type and tested across the other three. AVG denotes the average AUC score across three cross-manipulation trials. MoE-FFD outperforms the baseline in both intra- and cross-manipulation evaluations. Remarkably, MoE-FFD achieves around 99% AUC scores in each intra-manipulation evaluation, highlighting its outstanding detection accuracy in intra-domain evaluations. When compared to the state-of-the-art RECCE method, our MoE-FFD consistently achieves superior average cross-manipulation results across all four evaluation settings. These findings suggest that MoE-FFD is generalized to unseen manipulations.
TABLE IV: Ablation experiment on the designed MoE modules.
ViT-B MoE MoE CDF DFDC-P WDF
@@ 4.3 Comparison with State-of-the-Arts
4.3.3 Robustness to Real-World Perturbations.
Images and videos transmitted online always undergo various perturbations that erase forgery cues within the image/video contents [81]. As such, the detection performance of existing models significantly drops on distorted data. In this work, we introduce common perturbations to measure the model’s robustness, including Gaussian blurring, pink noise, white noise, and blockwise. Each perturbation type further involves five severity levels to mimic diverse real-world conditions. Note that we do not apply any data augmentations during training, such that the tested perturbations are totally unseen for our model. As shown in Fig. 4, the detectors’ AUC detection performance consistently deteriorates with the increasing severity level. However, our proposed MoE-FFD exhibits significantly greater resilience to most perturbations compared to previous methods. This indicates that the proposed MoE-FFD effectively captures the inherent forgery clues. Notably, MoE-FFD achieves substantial improvements in robustness compared to the ViT-B baseline, demonstrating that the designed MoE learning scheme and the PEFT modules greatly enhance the model’s robustness against image perturbations.
4.3.4 Discussion.
MoE-FFD achieves superior generalizability and robustness compared to previous methods, which can be attributed to the following designs: (1) MoE-FFD only updates external modules while preserving the abundant ImageNet knowledge, enabling the model to adaptively learn forgery-specific features; (2) MoE-FFD integrates the LoRA and Convpass Adapter with the ViT backbone, effectively leveraging the expressivity of transformers and the local forgery priors; (3) The incorporation of MoE modules facilitates optimal selection of LoRA and Adapter experts for forgery feature mining. Furthermore, MoE-FFD presents a parameter-efficient approach to face forgery detection due to the utilized PEFT strategy. To validate the effectiveness of the designed components, we perform ablation experiments next.
TABLE V: Cross-dataset detection performance on different ViT backbones.
Backbone MoE-FFD #Params CDF DFDC-P WDF
@@
4.4 Ablation Experiments
4.4.1 Impacts of the PEFT Layers.
The proposed MoE-FFD integrates one LoRA layer and one Adapter layer with each ViT block. The LoRA layer is designed to learn forgery-specific parameters for subsequent attention mechanisms, capturing long-range interactions within the input faces. Meanwhile, the Convpass Adapter introduces forgery local priors into the plain ViT model. To study the effectiveness of the designed PEFT modules, we report the cross-dataset detection results in Table IV. Utilizing MoE LoRA and MoE Adapter significantly improves the model generalizability across all datasets compared to the vanilla ViT-B backbone. This enhancement can be attributed to the LoRA and Adapter layers effectively retaining the ImageNet knowledge while adaptively learning forgery features. Furthermore, the combination of LoRA and Adapter layers allows the model to capture both long-range interactions and local forgery cues, further enhancing its face forgery detection performance.
4.4.2 Effectiveness on Other Backbones.
Flexibility is crucial in real-world applications to address the complexities of practical scenarios. Deployment devices with varying computational resources may necessitate different detection models. Therefore, we integrate the proposed MoE-FFD into ViT-Tiny, ViT-Small, and ViT-Large models to assess the flexibility of our approach. MoE-FFD can be readily inserted to other vision transformer backbones in a plug-and-play manner. The cross-dataset results are presented in Table V. Compared with directly finetuning the vanilla ViT backbones, MoE-FFD significantly reduces the training parameters. As the model size increases, the model generalizability consistently enhances. Additionally, significant performance improvements are observed across different ViT backbones by using MoE-FFD. Specifically, when we assemble MoE-FFD with the ViT-Large backbone, the model achieves 13.08%, 3.05%, and 7.04% AUC boosts on CDF, DFDC-P, and WDF datasets, with only $\sim$13.6trainable parameters. This, in turn, demonstrates the flexibility and effectiveness of our method.
TABLE VI: Impacts of the matrix rank of LoRA layers and the effectiveness of MoE.
Setting CDF DFDC-P WDF
@@ 4.4 Ablation Experiments
Figure 5: LoRA expert selection frequency on (a) CDF, (b) WDF, (c) DFDC-P, and (d) DFD datasets; Adapter expert selection frequency on (e) CDF, (f) WDF, (g) DFDC-P, and (h) DFD datasets.
4.4.3 Impacts of the MoE Learning Scheme.
In this work, the proposed MoE is designed to dynamically select the optimal Top-1 expert for face forgery detection. In Table VI, we compare the cross-dataset detection performance of the proposed MoE-FFD with individual LoRA experts of varying ranks. Notably, different datasets exhibit varying optimal LoRA ranks. The model reaches the best AUC scores on CDF, DFDC-P, and WDF datasets with the LoRA rank of 64, 96, and 48, respectively. MoE dynamically selects the LoRA rank, facilitating the model to search the optimal feature space for each input query. In Table VI, MoE consistently outperforms individual LoRA experts on all three datasets. Similarly, we investigate the impact of different Adapters in Table VII. As introduced in Sec. 3.3.2, the designed adapters tend to expose different local artifacts of input faces. MoE smartly searches the best local feature extractor for each input real/fake face, thereby achieving the superior results compared to using a single Adapter. It should be noted that the MoE approach only introduces negligible additional activated parameters, demonstrating that the performance improvements stem from the proposed MoE learning scheme, instead of the model scaling up.
In Fig. 5, we further investigate the expert selection distributions on four datasets. Fig. 5 (a)-(d) show the LoRA expert selection frequency on CDF, WDF, DFDC-P, and DFD datasets, while Fig. 5 (e)-(h) illustrate the Adapter expert selection frequency. Different Deepfake datasets generally exhibit significant domain gaps. From Fig. 5, we observe distinct LoRA and Adapter expert selection distributions among the four datasets. Our model dynamically projects the features into suitable space and extracts informative local features tailored to each dataset. This observation underscores MoE-FFD’s capability to select optimal experts tailored to different data, offering valuable insights into its adaptive nature.
TABLE VII: Impacts of different designed Adapters and the effectiveness of MoE.
Setting CDF DFDC-P WDF
@@
TABLE VIII: MoE v.s. Multi-Experts.
Method Training Inference CDF DFDC-P WDF
@@
While the proposed MoE learning scheme exhibits superior generalizability compared to using a single expert, it is still ambiguous whether the performance boosts stem from the designed MoE or the joint usage of experts. To further demonstrate the dynamic selection of optimal LoRA and Adapter experts by MoE, we compare MoE with Multi-Experts (Multi-E). Multi-E aggregates the features of all designed experts. The efficiency and the face forgery detection performance of MoE and Multi-E are reported in Table VIII. Thanks to the used gating mechanism that selectively activates the sparse experts, our MoE achieves a 1.28$\times$ speedup in training and 1.23$\times$ speedup in inference. Despite these efficiency gains, we interestingly find that MoE consistently outperforms Multi-E in terms of the cross-dataset detection performance. This suggests that naively aggregating multiple experts may not be the optimal strategy for face forgery detection. One potential explanation for this phenomenon is that using all feature extractor experts could suppress the most informative features while introducing noisy ones. The results in Table VIII further underscores the effectiveness of the proposed method in selecting the optimal expert for face forgery detection.
Figure 6: t-SNE feature distributions of the ViT baseline and MoE-FFD. Visualization results of (a) ViT baseline on CDF dataset, (b) MoE-FFD on CDF dataset, (c) ViT baseline on WDF dataset, and (d) MoE-FFD on WDF dataset.
Figure 7: Grad-CAM maps of the baseline model (ViT) and our proposed method MoE-FFD on six Deepfake datasets: FF++/DF, CDF, DFDC-P, WDF, DFD, and DFR.
@@
The gating network often exhibits a tendency to consistently assign large weights to only a few experts [48], resulting in overfitting problems. To address this issue, we introduce an MoE loss component aimed at encouraging equal importance among all experts. This regularization also prevents the model from getting trapped in local optima. In this subsection, we examine the impacts of the proposed $L_{moe}$. Table IX presents the cross-dataset evaluation performance with different values of the $L_{moe}$ loss weight $\lambda$ in Eq. (5), where $\lambda=0$ represents no MoE loss applied. We observe that the use of $L_{moe}$ effectively mitigates the gate overfitting problem and consistently boosts model generalizability. Furthermore, the model achieves the best generalizability with $\lambda=1$.
These findings highlight the importance of incorporating balanced expert contributions in the MoE framework. $L_{moe}$ component ensures that the learned representations are more diverse and generalizable across different datasets. This is particularly crucial in practical applications where data distribution can vary significantly. The optimal performance at $\lambda=1$ suggests that the balance between expert utilization and regularization is vital.
@@ 4.5 Visualization Results
We further provide the Grad-CAM maps of the baseline model and MoE-FFD in Fig. 7. Both models are trained on FF++ dataset and tested on six Deepfake datasets: FF++/DF, CDF, DFDC-P, WDF, DFD, and DFR datasets. While the baseline model often neglects informative fake facial regions or attends to peripheral irrelevant areas, MoE-FFD consistently directs attention to the manipulated regions within each input face. Despite the diverse environments captured in these datasets, including conditions like poor illumination, extreme head pose, and low resolution, MoE-FFD accurately localizes the forgery regions containing abundant forgery features. This observation further underscores the generalizability of our method.
TABLE IX: Effectiveness of loss components.
$L_{ce}$ $\lambda$ CDF DFDC-P WDF
AUC EER AUC EER AUC EER
0 82.65 25.57 78.14 29.06 76.29 30.43
0.1 85.09 23.29 80.05 26.96 79.56 27.92
1 86.69 22.06 80.83 26.67 80.64 27.11
5 84.01 24.69 81.47 26.68 80.14 27.32
10 83.35 25.53 80.46 28.13 78.93 29.05
5 Conclusions and Future Works
In this paper, we introduced MoE-FFD, a generalized yet parameter-efficient method for detecting face forgeries. By incorporating external lightweight LoRA and Adapter layers with the frozen ViT backbone, our framework adaptly acquired forgery-specific knowledge with minimal activated parameters. This approach not only harnesses the expressiveness of transformers but also capitalizes on the local forgery priors with customized adapters, contributing to enhanced detection performance. Through dynamic expert selection within both LoRA and Adapter layers, our MoE design further enhances the model’s generalizability and robustness. Extensive experiments consistently demonstrated MoE-FFD’s superiority in face forgery detection across diverse datasets, manipulation types, and perturbation scenarios. Moreover, MoE-FFD serves as a parameter-efficient detector and can seamlessly adapt to various ViT backbones, facilitating its deployment and fine-tuning in real-world applications. Last but not least, comprehensive ablation studies demonstrated the effectiveness of our designed LoRA layers and Convpass Adapter layers, and the MoE learning scheme indeed helped the model to search the optimal forgery experts. We anticipate that MoE-FFD will inspire future advancements in face forgery detection, particularly in bolstering generalization and efficiency.
While our MoE-FFD framework achieves generalized and robust performance in image-level face forgery detection, adapting our network design and learning scheme to video-level Deepfake detection is worth considering. Moreover, detecting audio-visual content forgery is essential in real world applications, multimodal Deepfake detection opens an important research path forward. We envision that our proposed MoE structure and the PEFT strategy should be still effective in these tasks. This path can be explored in our future work.